OpenStudy (anonymous):

How much will it cost to tile a kitchen floor that is 12 feet by 15 feet, with a circular island that has a diameter of 4 feet? Each case of tile costs $45.00 and can cover 15 square feet. Tiles are not sold individually. A) $135 B) $180 C) $495 D) $540

OpenStudy (anonymous):

quickly im taking a test!

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

its an assement test for my degree

OpenStudy (anonymous):

Of the fifty states in the United States, 12 have at least one mountain higher than 12,000 feet. What fraction of states do NOT have a mountain higher than 12,000 feet? A) - 6/25 B) 19/25 C) 5/1200 D) 1195/1200

OpenStudy (anonymous):

 area for tile = 180 - 4pi = 167 ft^2 167/15 = 11.16 = 12 12*45 D) $540 i think

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Given the following least squares prediction equation: y= -65x -173 we estimate y to_______ by_______ with each1-unit increase in x. A) Decrease, 65 B) Increase, 65 C) Decrease, 173 D) Increase, 173

OpenStudy (anonymous):

Of the fifty states in the United States, 12 have at least one mountain higher than 12,000 feet. What fraction of states do NOT have a mountain higher than 12,000 feet? A) - 6/25 B) 19/25 C) 5/1200 D) 1195/1200 38/50 = 19/25

OpenStudy (anonymous):

On a College Algebra test, the mean grade was 80% while the median grade was 84%. One student earned a grade of 83% and one student earned a grade of 85%. What can be inferred from this data? A) The majority of student grades were above 84%. B) More students earned grades of 84% than students who earned grades of 80%. C) At least one student earned a grade of 100%. D) The majority of student grades were above 80%.

OpenStudy (anonymous):

Given the following least squares prediction equation: y= -65x -173 we estimate y to_______ by_______ with each1-unit increase in x. A) Decrease, 65 B) Increase, 65 C) Decrease, 173 D) Increase, 173 A) Decrease, 65

OpenStudy (anonymous):

cheap one medal?

OpenStudy (anonymous):

On a College Algebra test, the mean grade was 80% while the median grade was 84%. One student earned a grade of 83% and one student earned a grade of 85%. What can be inferred from this data? B) More students earned grades of 84% than students who earned grades of 80%.